### v1

# An Enhanced Projection Pursuit Tree Classifier with Visual Methods for Assessing Algorithmic Improvements

Abstract
This paper presents enhancements to the projection pursuit tree classifier and visual diagnostic methods for assessing their impact in high dimensions. The original algorithm uses linear combinations of variables in a tree structure where depth is constrained to be less than the number of classes—a limitation that proves too rigid for complex classification problems. Our extensions improve performance in multi-class settings with unequal variance-covariance structures and nonlinear class separations by allowing more splits and more flexible class groupings in the projection pursuit computation. Proposing algorithmic improvements is straightforward; demonstrating their actual utility is not. We therefore develop two visual diagnostic approaches to verify that the enhancements perform as intended. Using high-dimensional visualization techniques, we examine model fits on benchmark datasets to assess whether the algorithm behaves as theorized. An interactive web application enables users to explore the behavior of both the original and enhanced classifiers under controlled scenarios. The enhancements are implemented in the R package PPtreeExt.
1 Introduction
Loh (2014) provides an extensive overview of decades of research on classification and regression tree algorithms. Tree models are useful because they are simple, yet flexible, fast to compute, widely applicable and interpretable. A main drawback, though, most available methods use splits along single variables, which means that separations best made using a combination of variables are awkward for trees to fit. While there have been many suggestions it is a much harder problem to optimize on linear combinations, and thus there is less available successful software for oblique splits.
One available method and software is the Projection Pursuit Tree (PPtree) algorithm (Y. D. Lee et al. 2013). It optimizes a projection pursuit index, for example the LDA (E.-K. Lee et al. 2005) or PDA (E.-K. Lee and Cook 2010) index among others, based on class information to identify one-dimensional projections that best separate the groups at each node. The PPtree tree structure is simpler than classic methods like rpart (Therneau and Atkinson 2025), as it restricts the depth to G1G-1 ($G$ is the number of classes). At each node, PPtree separates the data into two groups by first deciding which classes to group together; the data are split into left and right branches based on two group means. This produces a compact and interpretable tree implemented in PPtreeViz R package (E.-K. Lee 2018).
One of the main advantages of PPtree is its ability to leverage the correlation between predictor variables to improve class separation. It has been shown to outperform even random forests (RF) in some scenarios. Projection pursuit addresses a known limitation of the original RF algorithm, in which oblique projections were allowed. While it sounds appealing, the algorithm is practically infeasible because it is based on randomly generated splits. Thus it is slow, and inefficient, for finding useful separating oblique splits.
The compact nature of PPtree means that it fits simply, and does not require pruning. The resulting one-dimensional projections can be used to make visualizations of group separations. However, there are limitations that reduce its effectiveness in certain contexts. Due the the way classes are grouped, the prediction boundaries are often too close to one group, resulting in unnecessarily high error rates. This happens because grouping of classes likely produces unequal variance-covariances, making the pooled variance-covariance used by LDA inadequate. Moreover, when classes have non-elliptical variance-covariance, or are split into multiple clusters, the rigidity of the (G(G-1)-node structure, will not be sufficiently flexible.
A primary objective of this paper is to describe modifications to the PPtree algorithm that address these limitations. The algorithm is changed in two ways: (1) improving prediction boundaries by modifying the choice of split points-through class subsetting; and (2) increasing flexibility by allowing multiple splits per group. The modified algorithm is tested on a variety of benchmark data sets. These data sets include well-separated classes, the presence of outliers, and nonlinear boundaries. The performance is compared with the original PPtree algorithm and existing methods. Visual comparisons of decision boundaries, along with error rates, are used to further understand the model fits. The ultimate goal is to provide better building blocks for a projection pursuit-based classification random forest algorithm (da Silva, Cook, and Lee 2021) implemented in PPforest R package (da Silva, Cook, and Lee 2025).
Proposing enhancements to algorithms is easy, and theoretically the modifications may seem sensible. In practice, they may not work as one envisions. Thus, an important part of developing algorithm modifications is providing validation that they are doing what is expected. This means going beyond predictive accuracy, to understanding the resulting models. A second objective of this paper is to provide visual diagnostics for the algorithm changes. This is two-fold: a shiny (Chang et al. 2025) app and tours to view the fits in high dimensions. The app is available in the PPtreeExt package (da Silva, Cook, and Lee 2026) along with the implementation of the modified algorithm. It allows users to generate 2D data sets with different types of separations and compare the model fits for the modified and the original algorithm. This was an essential tool in constructing the new algorithm, because it provides side-by-side visualization of prediction boundaries to identify data patterns in which the different methods perform better. The second visual diagnostic is to use tours available in the tourr package (Wickham et al. 2011) to examine the high-dimensional structure of data sets and the resulting model fits where the modified algorithm outperforms other classifiers. This provides an explanation for the performance comparison results, and intuition for why the modified algorithm performs better.
The paper is organized as follows: Section 2 provides background on the original PPtree algorithm, including its structure and limitations. Section 3 describes the extensions to the algorithm. Section 4 presents the interactive visualization tool developed for exploring and comparing the algorithms. Section 5 presents the comparative analysis of algorithm performance, and visual diagnostics for the comparisons in Section 6. Finally, Section 7 discusses the changes, results of the comparison and potential future directions.
2 Background to PPtree
Figure 1 illustrates the original PPtree algorithm for $G=3$. Starting from the optimal projection, the green group-having the most distant mean—is relabeled as $g_{1}^{*}$, while the orange and violet groups are combined and relabeled as $g_{2}^{*}$. A second projection vector is then computed using all observations, following the same procedure applied to these two new super-groups. A split point is determined using one of the predefined rules, and observations are assigned to branches by comparing their projected values to the cutoff.
The algorithm computes a split point using one of several methods and the observations are allocated to groups based on whether their projected values fall below or above the cutoff. Since $g_{1}^{*}$ contains only one of the original classes, the left branch becomes a terminal node. Because $g_{2}^{*}$ still includes two original classes, the procedure is recursively repeated on the right branch.
Figure 1: Illustration of the original PPtree algorithm for $G=3$.
There are eight different rules available for computing the split point on a given projection. Let $\bar{\mathbf{x}}_{g}$, $\mathbf{x}_{g}^{med}$, $s_{g}$, $IQR_{g}$, and $n_{g}$ denote the mean, median, standard deviation, interquartile range, and sample size of group $g$, respectively. Table LABEL:tbl-rules summarizes the formulas used to calculate the split point value, $c$, under each rule.
Table 1: Rules for computing the split value, $c$, between two groups, on a data projection, using means, weighted mean, medians, standard deviations, interquartile ranges.
Rule Name Formula
1 mean $c=\frac{1}{2}\bar{\mathbf{x}}_{1}+\frac{1}{2}\bar{\mathbf{x}}_{2}$
2 sample size weighted mean $c=\left(\frac{n_{2}}{n_{1}+n_{2}}\right)\bar{\mathbf{x}}_{1}+\left(\frac{n_{1}}{n_{1}+n_{2}}\right)\bar{\mathbf{x}}_{2}$
3 standard deviation weighted mean $c=\left(\frac{s_{2}}{s_{1}+s_{2}}\right)\bar{\mathbf{x}}_{1}+\left(\frac{s_{1}}{s_{1}+s_{2}}\right)\bar{\mathbf{x}}_{2}$
4 standard error weighted mean $c=\frac{s_{2}/\sqrt{n_{2}}}{s_{1}/\sqrt{n_{1}}+s_{2}/\sqrt{n_{2}}}\bar{\mathbf{x}}_{1}+\frac{s_{1}/\sqrt{n_{1}}}{s_{1}/\sqrt{n_{1}}+s_{2}/\sqrt{n_{2}}}\bar{\mathbf{x}}_{2}$
5 median $c=\frac{1}{2}med_{1}+\frac{1}{2}med_{2}$
6 sample size weighted median $c=\frac{n_{2}}{n_{1}+n_{2}}med_{1}+\frac{n_{1}}{n_{1}+n_{2}}med_{2}$
7 IQR weighted median $c=\frac{IQR_{2}}{IQR_{1}+IQR_{2}}med_{1}+\frac{IQR_{1}}{IQR_{1}+IQR_{2}}med_{2}$
8 sample size and IQR weighted median $c=\frac{IQR_{2}/\sqrt{n_{2}}}{IQR_{1}/\sqrt{n_{1}}+IQR_{2}/\sqrt{n_{2}}}med_{1}+\frac{IQR_{1}/\sqrt{n_{1}}}{IQR_{1}/\sqrt{n_{1}}+IQR_{2}/\sqrt{n_{2}}}med_{2}$
2.1 Algorithm
The PPtree algorithm (Y. D. Lee et al. 2013) follows a multi-step procedure to identify linear combinations of predictors that optimally separate classes in multi-class settings. Let $d_{n}=\{(\mathbf{x}i,y_{i})\}_{i=1}^{n}$ denote the dataset, where $\mathbf{x}_{i}$ is a $p$-dimensional vector of explanatory variables. The corresponding class label is $y_{i}\in\mathscr{G}$, with $\mathscr{G}=\{1,2,\ldots,G\}$, for $i=1,\ldots,n$. The construction of a PPtree classifier follows a four-step algorithm, described as follows:
PPtree Algorithm (4-step version)
Input: Dataset $d_{n}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$, with $\mathbf{x}_{i}\in\mathbb{R}^{p}$, $y_{i}\in\mathscr{G}=\{1,\ldots,G\}$
Output: A projection pursuit classification tree with at most G1G-1 internal nodes
  • Step 1: Class Separation. Optimize a projection pursuit index (e.g., LDA or PDA) to find a projection vector $\boldsymbol{\alpha}$ that best separates all classes in the current node.
  • Step 2: Binary Relabeling. On projected data, $z_{i}=\boldsymbol{\alpha}^{T}\mathbf{x}_{i}$, compute projected class means $\bar{z}_{1},\ldots,\bar{z}_{G}$. Identify the pair with the largest mean difference and compute the midpoint $mp$. Relabel the observations into two super-classes, $g_{1}^{*}$ and $g_{2}^{*}$, depending on whether $\bar{z}_{g}$ is smaller than the midpoint $mp$.
  • Step 3: Node Splitting. Optimize a new projection $\boldsymbol{\alpha}^{*}$ to separate $g_{1}^{*}$ and $g_{2}^{*}$ based on the relabeled classes. Define the split value $c$ on $z_{i}^{*}=\boldsymbol{\alpha}^{*T}\mathbf{\bar{x}}_{gi^{*}}$ using a predefined rule (e.g., mean or weighted mean). Split the node using the rule $z_{i}^{*}<c$.
  • Step 4: Recursive Partitioning. Repeat Steps 1–3 for each child node until every terminal node contains only one original class.
2.2 Less desirable aspects
The decision boundaries of a classifier reflect how the model partitions the feature space and adapt to class separations. Visualizing the decision boundaries provides valuable insights into model behavior and limitations. To illustrate the motivation for extending the PPtree algorithm, we present simple examples based on simulated datasets. Details on the simulation settings are provided in Section 4.
Figure 2 shows the decision boundaries produced by rpart (left) and PPtree (right). In this example, the class structure is best separated using linear combinations of the two variables. The rpart algorithm, which relies on axis-aligned splits, must approximate the ideal boundary using multiple steps. In contrast, PPtree generates an oblique split that more effectively captures the separation, although the resulting boundary between the orange and purple classes lies too close to one of them.
The first split isolates the violet class from the other two. A second partition then attempts to separate the green and orange groups. Although the group structure is similar across the three classes, the resulting decision boundaries are not parallel. This occurs because the initial split uses the combined information from the orange and green groups-treated as a super-group—to compute the projection and determine the splitting point.
Alternative rules for computing the split value (see Table LABEL:tbl-rules) have a substantial impact on the position and orientation of these decision boundaries.
Figure 2: Comparison of decision boundaries produced by the rpart (left) and PPtree (right) algorithms on two-dimensional simulated data. The boundaries generated by PPtree are oblique to the coordinate axes, capturing the linear association between the two variables.
Figure 3: Comparison of decision boundaries produced by the rpart (left) and PPtree (right) algorithms on two-dimensional simulated data. The orange class cannot be separated using a single linear partition, and PPtree fails to model it accurately because each original class must be assigned to a single terminal node.
Figure 3 displays a simulated example that highlights the limited flexibility of the decision boundaries produced by PPtree. In this scenario, the orange class cannot be separated using a single linear partition, and PPtree fails to model it accurately because each original class must be assigned to a single terminal node. In contrast, the more traditional rpart algorithm performs better in this case, as its recursive structure allows it to accommodate the relatively simple non-linear separation.
3 PPtree extensions
The algorithm has been extended in two main ways: by modifying the rules used to determine decision boundaries and by allowing multiple splits per group. Two new approaches are introduced for computing the optimal split value. To support multiple splits within a group, additional stopping rules must be incorporated into the algorithm.
3.1 Modification 1: Subsetting classes to produce better boundaries
The first modification targets the third step of the original algorithm. Rather than combining all remaining classes into a single super-class, only the two closest ones—identified by the smallest difference in their projected means—are used to compute the next projection and define the split point
$3^{*}$ Find an optimal one-dimensional projection $\alpha^{**}$, using a subset of $\{(\mathbf{x_{i}},y_{i}^{*})\}_{i=1}^{n}$ to separate the two class problem $g_{1}^{*}$ and $g_{2}^{*}$. To this second projection only information from the two closest groups, one from $g_{1}^{*}$ and the other from $g_{2}^{*}$ are used. The best separation of $g_{1}^{*}$ and $g_{2}^{*}$ is determined by $\boldsymbol{\alpha}^{*T}\mathbf{\bar{x}}_{g1^{*}}<c$; then $g_{1}^{*}$ is assigned to the left node else assign $g_{2}^{*}$ to the right node, where $\mathbf{\bar{x}}_{g1^{*}}$ is the mean of $g_{1}^{*}$.
Figure 4 illustrates the modified version of the algorithm for a three-class classification problem. After identifying the initial optimal projection, instead of combining two groups into a single super-group, the group furthest from the others is temporarily excluded. In this example, class $g_{3}$ is dropped from the current node. A new projection is then computed using only classes $g_{1}$ and $g_{2}$ (relabeled as $g_{1}^{*}$ and $g_{2}^{*}$ ), and one of the eight available rules is applied to determine the optimal split point. Subsequently, class $g_{3}$ is reintroduced, and a new projection is computed to best separate it from $g_{2}$, completing the tree construction.
Figure 6 shows the effect of this modification on the decision boundaries obtained from the simulated dataset presented in Figure 2.
Figure 4: Illustration of the algorithm with Modification 1 applied to a three-class problem. At each node, the second one-dimensional projection is computed using only the two closest groups to determine the best projection direction and split point.
3.2 Modification 2: Multiple splits per class based on entropy reduction
This modification introduces a new approach for selecting the split value, based on the impurity of the resulting groups. This approach resembles the impurity-based criteria used in the rpart algorithm, but it operates on projected data rather than individual features. Currently, the only implemented impurity measure is negative entropy:
$E(s)=-\sum_{j=1}^{G}p_{js}\log(p_{js})$ (1)
where $p_{js}$ is the proportion of observations from class $j$ in subset $s$, and $G$ is the number of classes. Higher values of $E(s)$ indicate greater class impurity. When all observations in the subset belong to the same class, the entropy reaches its minimum value of $E(s)=0$, indicating a pure node. To evaluate the quality of a candidate split, the impurity values of the resulting left and right subsets are combined using a weighted average of the entropie:
$E(s_{L},s_{R})=\frac{n_{L}}{n_{L}+n_{R}}E(s_{L})+\frac{n_{R}}{n_{L}+n_{R}}E(s_{R})$ (2)
where $n_{L},n_{R}$ are the number of observations in the left and right subsets, respectively. The combined impurity, $E(s_{L},s_{R})$ is computed for all possible split points-defined as the midpoints between consecutive ordered projected values-denoted $\mathscr{C}=\{c_{1},...,c_{n_{L}+n_{R}-1}\}$. The split point that minimizes the combined entropy is selected:
${\min_{c\in\mathscr{C}}}E_{c}(s_{L},s_{R})$
This represents a major departure from the original PPtree structure, aiming to enhance flexibility in modeling non-linear class boundaries by allowing multiple splits within a class. Instead of relying on summary statistics at the class level, this approach shifts the focus to local differences between individual observations. To support this flexibility, additional control mechanisms are required, particularly new stopping criteria. While this is conceptually similar to the rpart algorithm, a key distinction is that larger sample sizes per node are needed to reliably estimate projection coefficients.
Figure 5 illustrates this modified version of the algorithm, which follows the steps outlined below.
PPtree Algorithm with Entropy-Based Splitting
Input: Dataset $d_{n}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$, with $\mathbf{x}_{i}\in\mathbb{R}^{p}$, $y_{i}\in\mathscr{G}=\{1,\ldots,G\}$
Output: A projection pursuit classification tree with entropy-based node splitting
  • Step 1: Class Separation. Optimize a projection pursuit index (e.g., LDA or PDA) to find a projection vector $\boldsymbol{\alpha}$ that best separates the classes in the current node.
  • Step 2: Projection and Candidate Splits. On the projected data, $z_{i}=\boldsymbol{\alpha}^{T}\mathbf{x}_{i}$, define candidate splits as midpoints between consecutive $z_{i}$ values.
  • Step 3: Impurity Evaluation. For each candidate split $c$, divide the data into $s_{L}=\{z_{i}<c\}$ and $s_{R}=\{z_{i}\geq c\}$. Compute the class proportions $p_{js}$ for each subset and evaluate the entropy:
$E(s)=-\sum_{j=1}^{G}p_{js}\log(p_{js})$
Compute the total impurity as:
$E(s_{L},s_{R})=\frac{n_{L}}{n_{L}+n_{R}}E(s_{L})+\frac{n_{R}}{n_{L}+n_{R}}E(s_{R})$
  • Step 4: Best Split Selection. Select the split point $c^{*}$ that minimizes $E(s_{L},s_{R})$. Split the node using the rule $z_{i}<c^{*}$.
  • Step 5: Recursive Partitioning. Repeat Steps 1–4 recursively on each child node until some of the following stopping criteria are met:
    • All observations in the node belong to the same class;
    • The node size is smaller than a minimum threshold $n_{s}$;
    • The entropy reduction is smaller than a threshold $ent_{s}$.
Figure 5: Illustration of the algorithm with Modification 2 applied to a three-class problem. Projections of the data are computed at each node, and multiple splits per class are allowed. An impurity-based criterion, such as entropy, is used to determine when and how often to split observations within a node.
4 Algorithm comparison
To evaluate both the original algorithm and the proposed modifications, we developed an interactive Shiny app (Chang et al. 2025). The tool includes methods for simulating 2D classification problems, and displays the decision boundaries generated by different algorithms: CART, PPtree, and the proposed extensions (implemented in rpart, PPtreeViz, and PPtreeExt, respectively). The Shiny app is available at https://natyds.shinyapps.io/explorapp/ and is implemented in PPtreeExt.
The app contains three tabs, each tab offers control over data generation and algorithm parameters: Basic-Sim, Sim-Outlier, and MixSim. The first tab simulates well-separated classes, the second includes a cluster of outliers, and the third generates overlapping groups using the MixSim package (Melnykov, Chen, and Maitra 2012). Each tab offers control over data generation and algorithm parameters. The decision boundaries for rpart, PPtree, and the modified PPtree are displayed side by side for visual comparison.
Additional controls, such as the rule partition used in PPtree algorithm and the modified PPtree based on subsetting clases, can be incorporated into the application. These features allow users to examine how changes in rule criteria influence the resulting classification boundaries.
Figure 6 illustrates the use of the interactive application. The first modification (Subsetting clases) is applied to generate the projection, and Rule 1 is used to compute the split point. The modified PPtree achieves an error rate comparable to the original PPtree, with both substantially outperforming rpart. Additionally, the resulting decision boundary more accurately partitions the groups in a balanced manner.
Figure 6: Model fits for the three algorithms—rpart, PPtree, and the modified PPtree based on subsetting clases applied to simulated well-separated classes. For PPtree, Rule 1 is used to determine the split points. The modified PPtree produces decision boundaries that are more symmetrically placed between the clusters.
Figure 7 presents a scenario in which Modification 2 (Multiple splits) yields a substantial improvement over the other algorithms. The use of multiple splits enables the small group of outliers from class 2 to be correctly classified.
Figure 7: Model fits for the three algorithms—rpart, PPtree, and the modified PPtree—applied to data in which one class consists of two distinct clusters. The boundaries produced by the modified PPtree with multiple splits better capture the underlying structure of the data compared to the other methods.
Figure 8 compares the performance of the algorithm variants on data simulated from Gaussian mixtures. The user can adjust several parameters of the mixture model: $n$ is the sample size, $BarOmega$ is the desired average overlap between components, $MaxOmega$ specifies the maximum allowable overlap, and $K$ is the number of groups.
Figure 8: Decision boundaries produced by the three algorithms— rpart, PPtree, and the modified PPtree—on data simulated from a Gaussian mixture model.
The interactive application provides insight into how modifications to the algorithm affect model fit. It follows the principle of visualizing the model in the data space (Wickham, Cook, and Hofmann 2015). An additional and important use of the tool has been for debugging and evaluating the algorithmic behavior. By adjusting the simulated data and model parameters, it becomes evident that small changes in the data can result in substantial changes in the fitted model—particularly when projections are involved.
This sensitivity is likely to be even greater in higher-dimensional settings. However, in scenarios where class separation occurs along combinations of variables, the use of projected data at each node tends to yield decision boundaries that more closely align with the true data distribution.
5 Performance comparison
This section presents a benchmark data study to evaluate the predictive performance of the proposed method. We compare the extended PPtree models with several classification methods—CART, PPtree, Random Forest, PPforest, and SVM—using 94 benchmark datasets. Table LABEL:tbl-source-tab1 in the Appendix contain the url with the source data while Table LABEL:tbl-data-tab2 summarize the main characteristics of each dataset summarizing the number of observations, predictors, number of classes, class imbalance, and absolute average correlation among predictors. Imbalance quantifies the disparity in class sizes (0 indicates perfectly balanced classes, biggest proportion size difference between classes), while higher correlation suggests a greater potential for class separation through linear combinations of variables.
For each dataset, two-thirds of the observations are randomly selected for training, and the remaining one-third is used for testing and computing the prediction error. This process is repeated 200 times.
Figure 9: Average test error from 200 training/test splits of benchmark data for four models: rpart, PPtree PPtreeExt modifications 1 and 2, where the PPtree extension performs better than rpart and the original PPtree.
Moreover, in 18 datasets, the modified PPtree outperforms CART, as illustrated in Figure 9, and in only 3 cases it achieves the lowest error among all selected methods, including ensemble-based classifiers. Table LABEL:tbl-datatbl presents an overview of these datasets, which is a subset of Table LABEL:tbl-data-tab2 from the Appendix.
Table 2: Overview of benchmark datasets where the proposed tree-based extensions outperform the selected tree methods. The table summarizes the number of observations, predictors, number of classes, class imbalance, and absolute average correlation among predictors. Imbalance quantifies the disparity in class sizes (0 indicates perfectly balanced classes), while higher correlation suggests a greater potential for class separation through linear combinations of variables.
dataset Cases Predictors Groups Imbalance Correlation
data20-3 8143 5 2 0.58 0.45
data27-a11 1747 18 5 0.38 0.24
data27-a31 1834 18 5 0.28 0.27
data27-b31 1424 18 5 0.21 0.24
data27-c11 1111 18 5 0.14 0.27
data27-c31 1448 18 5 0.17 0.27
data37-5 17560 53 5 0.21 0.16
data38 182 12 2 0.43 0.21
data63 6598 166 2 0.69 0.27
data34 1372 4 2 0.11 0.43
data49-2 606 100 2 0.03 1.00
data49-3 606 100 2 0.01 0.99
data22-1 258 5 4 0.25 0.12
data49-4 606 100 2 0.01 0.99
data56-1 7494 16 10 0.01 0.27
data67 20000 16 26 0.00 0.18
data69-1 5000 21 3 0.01 0.30
data69-2 5000 40 3 0.01 0.09
6 Explanation of performance
We would expect that the modified PPtree algorithm outperforms
  • rpart when the best separation between groups is in oblique directions
  • and the original PPtree when class variances are different, or when there are multiple classes.
This can be investigated visually using a tour. There are several strategies for approaching this:
  1. 1.
    On the data
    1. a.
      Work on two groups at a time, if there are more than 2
    2. b.
      Use a guided tour on the data to find the combinations of the variables where the groups are separated
  2. 2.
    Predict a grid of values that covers the data domain, for the methods interested in comparing, PPtreeExt, rpart and PPtree
    1. a.
      subset to the points near the border generated by the classifier.
    2. b.
      record a sequence of projection bases that shows the boundary created by PPtree
    3. c.
      use the same projection sequence to view the boundaries for the three different methods, possibly use a slice tour to focus on the boundary near the center of the data domain.
Figure 10 show projections of the data20-3 illustrating why PPtreeExt performs better. Group 0 is split on either side of group 1, which means that the original PPtree algorithm would not be able to capture both clusters. We can also see that the separation is mostly on V3 but small amounts of other variables help to improve the separation. This means that rpart will not work quite as well because it will focus on splitting V3.
(a) projection 1
(b) projection 2
Figure 10: Two projections from a grand tour of data20-3, from which we can see why the PPtreeExt method is the most effective. The group 0 is split into two clusters on either side of group 1, which means that the original algorithm is not optimal. Secondly, the separation between the two groups is in a combination of variables.
Figure 11 shows some plots illustrating why PPtreeExt outperforms the other methods of the data49-2 data. From Figure 9 we can see that the original PPtree (with LDA) algorithm performs much better than a tree, and the PPtreeExt performs much better again. This data has 100 predictors and 2 classes (Table LABEL:tbl-datatbl), which makes visualising more challenging that data20-3. A useful approach is to pre-process using principal component analysis (PCA) on the predictors, to reduce the number of variables to visualise, while maintaining class structure. Plot (a) shows PC 1 vs PC2, with points coloured by the two classes. PC 1 explains 99.8% of the variability in the predictors but it explains nothing of the class structure. This is not uncommon for PCA, where often interesting structure can be found in the smaller components. Plot (b) shows PC 2 vs PC 3, and here we see interesting class differences. Each class appears to have three linear pieces pointing in different directions. This supports why PPtreeExt might work well, because in the combination of variables, the classes are nearly separable with multiple linear cuts. It’s not feasible to continue working through principal components to tease apart the cluster structure relating to classes, so a good next step is to do projection pursuit. The optimal projection is shown in plot (c). It can be seen the two classes are mostly separable by a linear combination of variables, but there is no gap between the clusters. The linear combination with so many variables is what causes some difficulties for a simple tree algorithm. The interesting shapes of the clusters for each class, which appear to be multiple linear pieces, means that the PPtreeExt can find better separations than PPtree.
(a) PC1 vs PC2
(b) PC2 vs PC3
(c) optimal projection
Figure 11: Examining why data49-2 is best classified using PPtreeExt. Because there are 100 predictors, for visualisation it is useful to reduce the dimension using principal component analysis, and then examine the separation between groups. PC1 explains 99.8% of the overall variability among predictors, but it is not useful for separating the classes (a). PC2 and PC3 (b) have an interesting difference between the two classes which shows why multiple linear splits might be useful. A guided projection pursuit on the first 26 principal components (c) shows that these classes could be separated with a linear combination. This is why the original PPtree classifier does better than the tree algorithm and also why the extension improves the performance.
Figure 12 has plots illustrating why the PPtreeExt performs better on data69-1. From Figure 9 we see that all of the PPtree versions perform better than an rpart fit. This data has 21 predictors and 3 classes (Table LABEL:tbl-datatbl). This is sufficiently small to examine using a tour, and we initially use a guided tour with the LDA index. This produced the projection shown in plot (a) where we see the primary difference between the three classes. These are each similar size, but oriented in such a way that in a 2D projection they form a triangle. Plots (b) and (c) show the boundaries using a slice tour with the same projection as used in (a) for PPtreeExt and rpart, respectively. The boundary points are generated by simulating a random sample of observations between (-3, 3) in each dimension. The slice cuts through to centre of the box. We can see that PPtreeExt essentially fits the data shape neatly but the rpart boundary is very messy. This data needs oblique cuts, which PPtreeExt provides, and the modified algorithm handles the variance differences better than the original.
(a) data
(b) PPtreeExt
(c) rpart
Figure 12: Examining why data49-2 is best classified using PPtreeExt: (a) best projection from a guided tour shows that the data has three similar linear pieces oriented to form a triangle, (b) boundary produced by the PPtreeExt fitted model and (c) boundary produced by rpart fitted model.
7 Discussion
This article has presented two modifications to the PPtree algorithm for classification problems implemented in the R package PPtreeExt and evaluated using a purpose-built interactive web application and real data. The first modification involves subsetting classes when calculating the projection direction at each node, while the second allows for multiple splits per class based on individual entropy evaluations between candidate splits. These changes result in a more flexible classification method that leverages linear combinations of variables. The web application also demonstrates how easily diagnostic tools can be constructed to explore and understand the behavior of new algorithms.
The benchmark data study showed that PPtreeExt predictive performance is better than, CART and PPtree in 18 data out of 94. The modified PPtree algorithm outperforms CART in cases when the best separation between groups is in oblique directions and outperforms the original PPtree when class variances are different, or when there are multiple classes.
Interpretability is increasingly important, because predictions from algorithms may need to be defended in practice. We have shown how to examine the data and the fitted model to understand why the PPtree extension performs better than other tree methods on some data sets. This approach is generally useful for assessing and diagnosing model fits.
The modified tree algorithm can be incorporated into a projection pursuit forest classifier (da Silva, Cook, and Lee 2025). Given its implicit variable selection mechanism, the ensemble is expected to handle nuisance variables more effectively than a single tree. Additionally, the use of bagging should yield improved and more robust performance in problems with nonlinear decision boundaries. This can be accomplished using the current PPforest package.
8 Appendix
Table 3: Source data information
Name URL
data12 https://archive.ics.uci.edu/dataset/518/speaker+accent+recognition
data13-1 https://archive.ics.uci.edu/dataset/547/algerian+forest+fires+dataset
data13-2 https://archive.ics.uci.edu/dataset/547/algerian+forest+fires+dataset
data20-1 https://archive.ics.uci.edu/dataset/357/occupancy+detection
data20-2 https://archive.ics.uci.edu/dataset/357/occupancy+detection
data20-3 https://archive.ics.uci.edu/dataset/357/occupancy+detection
data21 https://archive.ics.uci.edu/dataset/537/cervical+cancer+behavior+risk
data22-1 https://archive.ics.uci.edu/dataset/257/user+knowledge+modeling
data22-2 https://archive.ics.uci.edu/dataset/257/user+knowledge+modeling
data26 https://archive.ics.uci.edu/dataset/308/gas+sensor+array+under+flow+modulation
data27-a11 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-a21 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-a31 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-a32 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-b11 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-b12 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-b31 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-b32 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-c11 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-c12 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-c31 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data27-c32 https://archive.ics.uci.edu/dataset/302/gesture+phase+segmentation
data29 https://archive.ics.uci.edu/dataset/282/lsvt+voice+rehabilitation
data32 https://archive.ics.uci.edu/dataset/257/user+knowledge+modeling
data34 https://archive.ics.uci.edu/dataset/267/banknote+authentication
data36 https://archive.ics.uci.edu/dataset/236/seeds
data37-1 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-2 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-3 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-4 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-5 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-8 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data37-14 https://archive.ics.uci.edu/dataset/231/pamap2+physical+activity+monitoring
data38 https://archive.ics.uci.edu/dataset/230/planning+relax
data39-1 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
Name URL
data39-2 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-3 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-4 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-5 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-6 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-7 https://archive.ics.uci.edu/dataset/224/gas+sensor+array+drift+dataset
data39-8 https://archive.ics.uci.edu/dataset/225/gas+sensor+array+drift+dataset
data39-9 https://archive.ics.uci.edu/dataset/226/gas+sensor+array+drift+dataset
data41-1 https://archive.ics.uci.edu/dataset/194/wall+following+robot+navigation+data
data41-2 https://archive.ics.uci.edu/dataset/194/wall+following+robot+navigation+data
data41-3 https://archive.ics.uci.edu/dataset/194/wall+following+robot+navigation+data
data42 https://archive.ics.uci.edu/dataset/192/breast+tissue
data44 https://archive.ics.uci.edu/dataset/181/libras+movement
data49-2 https://archive.ics.uci.edu/dataset/167/hill+valley
data49-3 https://archive.ics.uci.edu/dataset/168/hill+valley
data49-4 https://archive.ics.uci.edu/dataset/169/hill+valley
data50 https://archive.ics.uci.edu/dataset/161/mammographic+mass
data52-1 https://archive.ics.uci.edu/dataset/96/spectf+heart
data52-2 https://archive.ics.uci.edu/dataset/96/spectf+heart
data53 https://archive.ics.uci.edu/dataset/139/synthetic+control+chart+time+series
data55-1 https://archive.ics.uci.edu/dataset/80/optical+recognition+of+handwritten+digits
data55-2 https://archive.ics.uci.edu/dataset/80/optical+recognition+of+handwritten+digits
data56-1 https://archive.ics.uci.edu/dataset/81/pen+based+recognition+of+handwritten+digits
data56-2 https://archive.ics.uci.edu/dataset/81/pen+based+recognition+of+handwritten+digits
data57 https://archive.ics.uci.edu/dataset/39/ecoli
data61-1 https://archive.ics.uci.edu/dataset/54/isolet
data62 https://archive.ics.uci.edu/dataset/74/musk+version+1
data63 https://archive.ics.uci.edu/dataset/75/musk+version+2
data64-1 https://archive.ics.uci.edu/dataset/146/statlog+landsat+satellite
data64-2 https://archive.ics.uci.edu/dataset/146/statlog+landsat+satellite
data65 https://archive.ics.uci.edu/dataset/15/breast+cancer+wisconsin+original
data66 https://archive.ics.uci.edu/dataset/62/lung+cancer
data67 https://archive.ics.uci.edu/dataset/59/letter+recognition
data68-1 https://archive.ics.uci.edu/dataset/50/image+segmentation
data68-2 https://archive.ics.uci.edu/dataset/50/image+segmentation
data69-1 https://archive.ics.uci.edu/dataset/107/waveform+database+generator+version+1
data69-2 https://archive.ics.uci.edu/dataset/107/waveform+database+generator+version+1
data70 https://archive.ics.uci.edu/dataset/42/glass+identification
data71-1 https://archive.ics.uci.edu/dataset/148/statlog+shuttle
Name URL
data71-2 https://archive.ics.uci.edu/dataset/148/statlog+shuttle
data72-1 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-2 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-3 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-4 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-5 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-6 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data72-7 https://archive.ics.uci.edu/dataset/149/statlog+vehicle+silhouettes
data73 https://archive.ics.uci.edu/dataset/151/connectionist+bench+sonar+mines+vs+rocks
data74 https://archive.ics.uci.edu/dataset/152/connectionist+bench+vowel+recognition+deterding+data
NCI60 https://github.com/natydasilva/PPforest/blob/master/data/NCI60.rda
crab https://github.com/natydasilva/PPforest/blob/master/data/crab.rda
fishcatch https://github.com/natydasilva/PPforest/blob/master/data/fishcatch.rda
glass https://github.com/natydasilva/PPforest/blob/master/data/glass.rda
image https://github.com/natydasilva/PPforest/blob/master/data/image.rda
leukemia https://github.com/natydasilva/PPforest/blob/master/data/image.rda
lymphoma https://github.com/natydasilva/PPforest/blob/master/data/lymphoma.rda
olive https://github.com/natydasilva/PPforest/blob/master/data/olive.rda
parkinson https://github.com/natydasilva/PPforest/blob/master/data/parkinson.rda
wine https://github.com/natydasilva/PPforest/blob/master/data/wine.rda
Table 6: Overview of benchmark datasets
dataset Cases Predictors Groups Imbalance Correlation
data12 329 12 6 0.41 0.38
data13-1 122 13 2 0.03 0.41
data13-2 121 13 2 0.29 0.37
data20-1 2665 5 2 0.27 0.81
data20-2 9752 5 2 0.58 0.29
data20-3 8143 5 2 0.58 0.45
data21 72 19 2 0.42 0.22
data22-1 258 5 4 0.25 0.12
data22-2 145 5 4 0.14 0.18
data26 58 435 4 0.21 0.64
data27-a11 1747 18 5 0.38 0.24
data27-a21 1264 18 5 0.35 0.22
data27-a31 1834 18 5 0.28 0.27
data27-a32 1830 32 5 0.28 0.13
data27-b11 1072 18 5 0.31 0.27
data27-b12 1069 32 5 0.32 0.13
data27-b31 1424 18 5 0.21 0.24
data27-b32 1420 32 5 0.21 0.11
data27-c11 1111 18 5 0.14 0.27
data27-c12 1107 32 5 0.13 0.11
data27-c31 1448 18 5 0.17 0.27
data27-c32 1444 32 5 0.17 0.11
data29 126 310 2 0.33 0.28
data32 403 6 4 0.20 0.11
data34 1372 4 2 0.11 0.43
data36 210 7 3 0.00 0.61
data37-1 29101 53 5 0.23 0.17
data37-14 769 53 2 0.52 0.16
data37-2 14132 53 3 0.62 0.18
data37-3 11835 53 4 0.34 0.16
data37-4 16357 53 5 0.29 0.15
data37-5 17560 53 5 0.21 0.16
dataset Cases Predictors Groups Imbalance Correlation
data37-8 22990 53 9 0.27 0.17
data38 182 12 2 0.43 0.21
data39-1 445 128 6 0.15 0.55
data39-2 1239 128 5 0.35 0.40
data39-3 1586 128 5 0.17 0.56
data39-4 161 128 5 0.32 0.43
data39-5 197 128 5 0.22 0.57
data39-6 2300 128 6 0.25 0.61
data39-7 3613 128 6 0.11 0.68
data39-8 294 128 6 0.43 0.57
data39-9 470 128 6 0.10 0.42
data41-1 5456 2 4 0.34 0.03
data41-2 5456 4 4 0.34 0.15
data41-3 5456 24 4 0.34 0.17
data42 106 9 6 0.08 0.51
data44 360 90 15 0.00 0.27
data49-2 606 100 2 0.03 1.00
data49-3 606 100 2 0.01 0.99
data49-4 606 100 2 0.01 0.99
data50 830 5 2 0.03 0.23
data52-1 80 44 2 0.00 0.25
data52-2 187 44 2 0.84 0.29
data53 600 60 6 0.00 0.46
data55-1 3823 64 10 0.00 0.12
data55-2 1797 64 10 0.01 0.12
data56-1 7494 16 10 0.01 0.27
data56-2 3498 16 10 0.01 0.28
data57 327 7 5 0.38 0.23
data61-1 6238 617 26 0.00 0.17
data62 476 166 2 0.13 0.26
data63 6598 166 2 0.69 0.27
data64-1 4435 36 6 0.15 0.49
data64-2 2000 36 6 0.13 0.50
data65 683 9 2 0.30 0.60
data66 27 56 3 0.07 0.20
dataset Cases Predictors Groups Imbalance Correlation
data67 20000 16 26 0.00 0.18
data68-1 210 19 7 0.00 0.30
data68-2 2100 19 7 0.00 0.28
data69-1 5000 21 3 0.01 0.30
data69-2 5000 40 3 0.01 0.09
data70 214 9 6 0.31 0.23
data71-1 43500 9 7 0.78 0.19
data71-2 14494 9 5 0.79 0.20
data72-1 94 18 4 0.09 0.39
data72-2 94 18 4 0.05 0.40
data72-3 94 18 4 0.10 0.42
data72-4 94 18 4 0.12 0.42
data72-5 94 18 4 0.10 0.49
data72-6 94 18 4 0.06 0.40
data72-7 94 18 4 0.06 0.44
data73 208 60 2 0.07 0.23
data74 990 12 11 0.00 0.22
NCI60 56 31 7 0.05 0.56
crab 200 6 4 0.00 0.95
fishcatch 159 7 7 0.31 0.46
glass 214 10 6 0.31 0.23
image 2310 19 7 0.00 0.28
leukemia 72 41 3 0.40 0.44
lymphoma 80 51 3 0.41 0.75
olive 572 9 9 0.32 0.35
parkinson 195 23 2 0.51 0.50
wine 178 14 3 0.13 0.30