### v2

# An Enhanced Projection Pursuit Tree Classifier with Visual Methods for Assessing Algorithmic Improvements

Abstract
This paper presents enhancements to the projection pursuit tree classifier and visual diagnostic methods for assessing their impact in high dimensions. The original algorithm uses linear combinations of variables in a tree structure where depth is constrained to be less than the number of classes—a limitation that proves too rigid for complex classification problems. Our extensions improve performance in multi-class settings with unequal variance-covariance structures and nonlinear class separations by allowing more splits and more flexible class groupings in the projection pursuit computation. Proposing algorithmic improvements is straightforward; demonstrating their actual utility is not. We therefore develop two visual diagnostic approaches to verify that the enhancements perform as intended. Using high-dimensional visualization techniques, we examine model fits on benchmark datasets to assess whether the algorithm behaves as theorized. An interactive web application enables users to explore the behavior of both the original and enhanced classifiers under controlled scenarios. The enhancements are implemented in the R package PPtreeExt, which is available on CRAN.
Natalia da Silva1   Dianne Cook2   Eun-Kyung Lee3
1Instituto de Estadística (IESTA), Universidad de la República, Montevideo, Uruguay
2Department of Econometrics and Business Statistics, Monash University, Melbourne, Australia
3Department of Statistics, Ewha Womans University, Seoul, Republic of Korea
Keywords: classification, heterogeneous variance, machine learning, projection pursuit, separability, tree classifier
1 Introduction
Loh (2014) provides an extensive overview of decades of research on classification and regression tree algorithms. Tree models are useful because they are simple, yet flexible, fast to compute, widely applicable, and interpretable. A main drawback, though, is that most available methods use splits along single variables, making separations on a combination of variables harder to capture. The reason is that it is a much harder problem to optimize the search for linear combinations, and thus, there is a lack of useful software that provides oblique splits.
One available method and software is the projection pursuit tree (PPtree) (Lee et al. 2013). It optimizes a projection pursuit (PP) index, for example, the LDA (Lee et al. 2005) or PDA (Lee and Cook 2010) index among others, based on class information to identify one-dimensional projections that best separate the groups at each node. The PPtree tree structure is simpler than classic methods like rpart (Therneau and Atkinson 2025), as it restricts the tree depth to G1G-1 ($G$ is the number of classes). At each node, PPtree separates the data into two groups by first deciding which classes to group together; the data are split into left and right branches based on two group means. This produces a compact and interpretable tree implemented in the PPtreeViz R package (Lee 2018).
One of the main advantages of PPtree is its ability to leverage the correlation between predictor variables to improve class separation. In some scenarios, it has been shown to outperform other classifiers, including random forests (RF) (Breiman 2001). PP addresses a known limitation of the original RF algorithm. It included oblique projections, but because it is based on randomly generated splits was practically infeasible. It is an inefficient algorithm for finding useful oblique splits.
The compact nature of PPtree means that it fits simply and does not require pruning. The resulting one-dimensional projections can be used to make visualizations of group separations. However, there are limitations that reduce its effectiveness. In multiclass problems, due to the way classes are grouped, the prediction boundaries are often too close to one group, resulting in unnecessarily high error rates. This happens because grouping of classes likely produces unequal variance-covariances, making the pooled variance-covariance used by LDA inadequate. Secondly, when classes have non-elliptical variance-covariance, or have multiple clusters per class, the rigidity of the (G(G-1)-node structure is not sufficiently flexible. The algorithm is modified to address these limitations in two ways: (1) improving prediction boundaries by modifying the choice of split points through class subsetting; and (2) increasing flexibility by allowing multiple splits per group.
Proposing enhancements to algorithms is easy, and theoretically, the modifications may seem sensible. In practice, they may not work as one envisions. Thus, an important part of developing algorithm modifications is providing validation that they are doing what is expected, or not. This means going beyond predictive accuracy to understanding the resulting models. A second objective of this paper is to provide visual diagnostics for the algorithm changes. This is two-fold: a shiny (Chang et al. 2025) app and tours to view the fits in high dimensions. The app is available in the PPtreeExt package (da Silva et al. 2026) along with the implementation of the modified algorithm. It allows users to generate 2D data sets with different types of separations and compare the model fits for the modified and the original algorithm. This was an essential tool in constructing the new algorithm, because it provides a side-by-side visualization of prediction boundaries to identify data patterns in which the different methods perform better. The second visual diagnostic is to use tours available in the tourr package (Wickham et al. 2011) to examine the high-dimensional structure of data sets and the resulting model fits, where the modified algorithm outperforms other classifiers. This explains the performance comparison results and intuition for why the modified algorithm performs better.
The paper is organized as follows: Section 2 provides background on the original PPtree algorithm, including its structure and limitations. Section 3 describes the extensions to the algorithm. Section 4 contains the examples on visually and numerically digesting the enhanced algorithm’s performance. Finally, Section 5 discusses the changes, results of the comparison, and potential future directions.
2 Background to PPtree
Figure 1 illustrates the original PPtree algorithm for $G=3$. Starting from the optimal projection, the left-side, most distant, cluster (green circles) is relabeled as $g_{1}^{*}$, while the other two clusters are combined and relabeled as $g_{2}^{*}$. A second projection vector is then computed using all observations, following the same procedure applied to these two new super-groups. A split point is determined using one of the predefined rules, and observations are assigned to branches by comparing their projected values to the cutoff.
The algorithm computes a split point using one of several methods, and the observations are allocated to groups based on whether their projected values fall below or above the cutoff. Since $g_{1}^{*}$ contains only one of the original classes, the left branch becomes a terminal node. Because $g_{2}^{*}$ still includes two original classes, the procedure is recursively repeated on the right branch.
Figure 1: Illustration of the original PPtree algorithm for $G=3$. In step 1, $P$ optimization decides on the best linear projection. Step 2 combines two groups to make a two-group classification. The $P$ optimization is computed for two group problem, and an appropriate split is decided. This is repeated for the remaining two groups to get the final model containing two splits. The locations where modifications will be made are circled.
There are eight different rules available for computing the split point on a given projection. Let $\bar{\mathbf{x}}_{g}$, $\mathbf{x}_{g}^{med}$, $s_{g}$, $IQR_{g}$, and $n_{g}$ denote the mean, median, standard deviation, interquartile range, and sample size of group $g$, respectively. Table LABEL:tbl-rules summarizes the formulas used to calculate the split point value, $c$, under each rule.
Table 1: Rules for computing the split value, $c$, between two groups, on a data projection, using means, weighted mean, medians, standard deviations, and interquartile ranges.
Rule Name Formula
1 mean $c=\frac{1}{2}\bar{\mathbf{x}}_{1}+\frac{1}{2}\bar{\mathbf{x}}_{2}$
2 sample size weighted mean $c=\left(\frac{n_{2}}{n_{1}+n_{2}}\right)\bar{\mathbf{x}}_{1}+\left(\frac{n_{1}}{n_{1}+n_{2}}\right)\bar{\mathbf{x}}_{2}$
3 standard deviation weighted mean $c=\left(\frac{s_{2}}{s_{1}+s_{2}}\right)\bar{\mathbf{x}}_{1}+\left(\frac{s_{1}}{s_{1}+s_{2}}\right)\bar{\mathbf{x}}_{2}$
4 standard error weighted mean $c=\frac{s_{2}/\sqrt{n_{2}}}{s_{1}/\sqrt{n_{1}}+s_{2}/\sqrt{n_{2}}}\bar{\mathbf{x}}_{1}+\frac{s_{1}/\sqrt{n_{1}}}{s_{1}/\sqrt{n_{1}}+s_{2}/\sqrt{n_{2}}}\bar{\mathbf{x}}_{2}$
5 median $c=\frac{1}{2}med_{1}+\frac{1}{2}med_{2}$
6 sample size weighted median $c=\frac{n_{2}}{n_{1}+n_{2}}med_{1}+\frac{n_{1}}{n_{1}+n_{2}}med_{2}$
7 IQR weighted median $c=\frac{IQR_{2}}{IQR_{1}+IQR_{2}}med_{1}+\frac{IQR_{1}}{IQR_{1}+IQR_{2}}med_{2}$
8 sample size and IQR weighted median $c=\frac{IQR_{2}/\sqrt{n_{2}}}{IQR_{1}/\sqrt{n_{1}}+IQR_{2}/\sqrt{n_{2}}}med_{1}+\frac{IQR_{1}/\sqrt{n_{1}}}{IQR_{1}/\sqrt{n_{1}}+IQR_{2}/\sqrt{n_{2}}}med_{2}$
2.1 Algorithm
The PPtree algorithm (Lee et al. 2013) follows a multi-step procedure to identify linear combinations of predictors that optimally separate classes in multi-class settings. Let $d_{n}=\{(\mathbf{x}i,y_{i})\}_{i=1}^{n}$ denote the dataset, where $\mathbf{x}_{i}$ is a $p$-dimensional vector of explanatory variables. The corresponding class label is $y_{i}\in\mathscr{G}$, with $\mathscr{G}=\{1,2,\ldots,G\}$, for $i=1,\ldots,n$. The construction of a PPtree classifier follows a four-step algorithm, as follows:
Input: Dataset $d_{n}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$, with $\mathbf{x}_{i}\in\mathbb{R}^{p}$, $y_{i}\in\mathscr{G}=\{1,\ldots,G\}$
Output: A projection pursuit classification tree with at most G1G-1 internal nodes
  • Step 1: Class Separation. Optimize a PP index (e.g., LDA or PDA) to find a projection vector $\boldsymbol{\alpha}$ that best separates all classes in the current node.
  • Step 2: Binary Relabeling. On projected data, $z_{i}=\boldsymbol{\alpha}^{T}\mathbf{x}_{i}$, compute projected class means $\bar{z}_{1},\ldots,\bar{z}_{G}$. Identify the pair with the largest mean difference and compute the midpoint $mp$. Relabel the observations into two super-classes, $g_{1}^{*}$ and $g_{2}^{*}$, depending on whether $\bar{z}_{g}$ is smaller than the midpoint $mp$.
  • Step 3: Node Splitting. Optimize a new projection $\boldsymbol{\alpha}^{*}$ to separate $g_{1}^{*}$ and $g_{2}^{*}$ based on the relabeled classes. Define the split value $c$ on $z_{i}^{*}=\boldsymbol{\alpha}^{*T}\mathbf{\bar{x}}_{gi^{*}}$ using a predefined rule (e.g., mean or weighted mean). Split the node using the rule $z_{i}^{*}<c$.
  • Step 4: Recursive Partitioning. Repeat Steps 1–3 for each child node until every terminal node contains only one original class.
2.2 Less desirable aspects
The decision boundaries of a classifier reflect how the model partitions the feature space and adapts to class separations. Visualizing the decision boundaries provides valuable insights into model behavior and limitations. To illustrate the PPtree algorithm relative to the classical rpart and the motivation for improving PPtree, we use two simulated three-class 2D data sets. Data set LC has separations on linear combinations of variables (Figure 2), and data set MC has multiple clusters in a class (Figure 3). The three classes are indicated by squares (S, violet), circles (C, green), and triangles (T, orange).
Figure 2 shows the decision boundaries produced by rpart (left) and PPtree (right), and percentage error, for the LC data. The class clusters are best separated using linear combinations of the two variables. The rpart algorithm, which relies on axis-aligned splits, must approximate the ideal boundary using rectangular steps. Even with many steps, it cannot be as efficient as the oblique splits provided by PPtree. The two steps of PPtree are to split the S cluster (violet) first, and then split C (green) and T (orange). Although the group structure is similar across the three classes, the resulting decision boundaries are not parallel. This occurs because the initial split uses the combined information from the C and T clusters - treated as a super class - to compute the projection and determine the splitting point. This also results in the boundary between T and S being too close to T. Alternative rules for computing the split value (see Table LABEL:tbl-rules) can impact the position and orientation of these decision boundaries, but in this case does not fix the deficiency.
Figure 2: Comparison of decision boundaries produced by the rpart (left) and PPtree (right) algorithms on the LC data. The boundaries generated by PPtree are oblique to the coordinate axes, appropriately capturing the separations as linear combinations of the two variables. Even with many splits, the rpart algorithm cannot elegantly classify this data.
Figure 3: Comparison of decision boundaries produced by the rpart (left) and PPtree (right) algorithms on the MC data. The T class (orange triangles) cannot be separated using a single linear partition, and PPtree fails to model it accurately because each class must be assigned to a single terminal node. This illustrates the need for allowing more splits in PPtree to capture unusually shaped clusters and separations involving combinations of variables.
Figure 3 compares the rpart and PPtree decision boundaries for the data set MC, where class T has two separated clusters. It highlights the limited flexibility of the decision boundaries produced by PPtree. In this scenario, the T cannot be separated using a single linear partition, and PPtree fails to capture it because the tree depth is limited to a single terminal node for each class. In contrast, the more traditional rpart algorithm performs better in this case, as its recursive structure allows it to accommodate the relatively simple non-linear separation. However, rpart fails to capture the separation between classes as linear combinations. This shows why more splits should be allowed for the PPtree algorithm.
3 PPtree extensions
The algorithm has been extended in two main ways, as indicated by the annotation in Figure 1. Section 4.1 explains how we can compare the changes these modifications induce on the boundaries relative to rpart and PPtree using the Shiny app.
3.1 Modification 1: Subsetting classes to produce better boundaries
The first modification targets the re-grouping step of the original algorithm. Rather than combining all remaining classes into a single super-class, only the two closest ones, as identified by the smallest difference in their projected means, are used to compute the next projection and define the split point. This step is now as follows:
  • Step 2: Binary Relabeling. Find an optimal one-dimensional projection $\alpha^{**}$, using a subset of $\{(\mathbf{x_{i}},y_{i}^{*})\}_{i=1}^{n}$ to separate the two class problem $g_{1}^{*}$ and $g_{2}^{*}$. To this second projection only information from the two closest groups, one from $g_{1}^{*}$ and the other from $g_{2}^{*}$ are used. The best separation of $g_{1}^{*}$ and $g_{2}^{*}$ is determined by $\boldsymbol{\alpha}^{*T}\mathbf{\bar{x}}_{g1^{*}}<c$; then $g_{1}^{*}$ is assigned to the left node else assign $g_{2}^{*}$ to the right node, where $\mathbf{\bar{x}}_{g1^{*}}$ is the mean of $g_{1}^{*}$.
In Figure 1 this would correspond to dropping class $g_{3}$ from the current node. The projection is then computed using only classes $g_{1}$ and $g_{2}$ (relabeled as $g_{1}^{*}$ and $g_{2}^{*}$ ). The choice of the existing eight available rules is then applied to determine the optimal split point, to create the terminal node for $g_{1}$. Subsequently, class $g_{3}$ is reintroduced, and a new projection is computed to best separate it from $g_{2}$. The split is made on this set of observations, creating two more terminal nodes, completing the tree construction.
3.2 Modification 2: Multiple splits per class based on entropy reduction
To allow for multiple splits, various changes are needed: the split is made on the optimal projection, not after groups are combined, and new stopping rules are defined to prevent over-fitting. Figure 4 illustrates the change to the algorithm for this modification.
In PPtree, the decision boundary split is made using the means for the two groups, just like is done with linear discriminant analysis. This modification changes this to follow the rpart algorithm, minimizing entropy, with the difference being that it is done on the optimal 1-D projected data. That is, the modification computes the negative entropy criteria:
$E(s)=-\sum_{j=1}^{G}p_{js}\log(p_{js})$ (1)
where $p_{js}$ is the proportion of observations from class $j$ in subset $s$, and $G$ is the number of classes. Lower values of $E(s)$ indicate purity, reaching zero when all observations in the subset belong to the same class. The impurity value for a split is calculated by combining the values of the left and right subsets:
$E(s_{L},s_{R})=\frac{n_{L}}{n_{L}+n_{R}}E(s_{L})+\frac{n_{R}}{n_{L}+n_{R}}E(s_{R})$ (2)
where $n_{L},n_{R}$ are the number of observations in the left and right subsets, respectively. Across all possible splits, the lowest combined impurity, $E(s_{L},s_{R})$, is computed for all possible split points, defined as the midpoints between consecutive ordered projected values, denoted $\mathscr{C}=\{c_{1},...,c_{n_{L}+n_{R}-1}\}$. The split point that minimizes the combined entropy is selected:
${\min_{c\in\mathscr{C}}}E_{c}(s_{L},s_{R})$
In addition, three stopping rules are introduced, which are similar to rpart: the node is pure, there are fewer than a specified number of observations in the node, or the reduction in impurity is less than a specified tolerance value. These changes provide the modified algorithm to be:
Input: Dataset $d_{n}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{n}$, with $\mathbf{x}_{i}\in\mathbb{R}^{p}$, $y_{i}\in\mathscr{G}=\{1,\ldots,G\}$
Output: A projection pursuit classification tree with entropy-based node splitting
  • Step 1: Class Separation. Optimize a PP index (e.g., LDA or PDA) to find a projection vector $\boldsymbol{\alpha}$ that best separates the classes in the current node.
  • Step 2: Projection and Candidate Splits. On the projected data, $z_{i}=\boldsymbol{\alpha}^{T}\mathbf{x}_{i}$, define candidate splits as midpoints between consecutive $z_{i}$ values.
  • Step 3: Impurity Evaluation. For each candidate split $c$, divide the data into $s_{L}=\{z_{i}<c\}$ and $s_{R}=\{z_{i}\geq c\}$. Compute the class proportions $p_{js}$ for each subset and evaluate the entropy:
$E(s)=-\sum_{j=1}^{G}p_{js}\log(p_{js})$
Compute the total impurity as:
$E(s_{L},s_{R})=\frac{n_{L}}{n_{L}+n_{R}}E(s_{L})+\frac{n_{R}}{n_{L}+n_{R}}E(s_{R})$
  • Step 4: Best Split Selection. Select the split point $c^{*}$ that minimizes $E(s_{L},s_{R})$. Split the node using the rule $z_{i}<c^{*}$.
  • Step 5: Recursive Partitioning. Repeat Steps 1–4 recursively on each child node until some of the following stopping criteria are met:
    • All observations in the node belong to the same class;
    • The node size is smaller than a minimum threshold $n_{s}$;
    • The entropy reduction is smaller than a threshold $ent_{s}$.
Figure 4: Illustration of the algorithm with modification 2 applied to a three-class problem. Projections of the data are computed at each node, and multiple splits per class are allowed. An impurity-based criterion, such as entropy, is used to determine when and how often to split observations within a node.
4 Comparative performance
Three approaches are used to examine the effectiveness of the modifications. The first is a Shiny app which generates structures in 2-D and shows the model fits for three algorithms: rpart, PPtree, and the PPtree extensions. It was developed and used to decide on effective changes to the algorithm. The second is a simulation study on 94 benchmark data sets to compare test errors. Lastly, to understand where, why, and how the extensions work, we examine the data visually in high dimensions along with the fitted models.
4.1 Exploring the algorithm changes interactively
The Shiny app includes methods for simulating 2D classification problems, and displays the decision boundaries generated by three different algorithms: CART, PPtree, and the proposed extensions as implemented in the rpart, PPtreeViz, and PPtreeExt packages, respectively.
There are three data generation methods, “Basic-Sim”, “Sim-Outlier”, and “MixSim”. Basic-Sim simulates classes with separations between clusters on linear combinations of variables, by user-specified correlations between variables for each cluster. Sim-Outlier uses the same approach with the addition of a cluster of outliers corresponding to one of the classes. These two methods allow only for three-cluster data. The user can control the number of observations in each class, the class means, and the Sim-Outlier mean and shape of the outlier groups. MixSim uses a Gaussian mixture model using the MixSim package (Melnykov et al. 2012) to simulate class clusters, with any number of classes. While all classes have equal sample size, the user can choose this and also specify values for the MixSim algorithm. Figure 5, Figure 6, Figure 7 show the interface for the three different data-generating methods. Clicking “Ok” will simulate new samples.
The decision boundaries for rpart, PPtree, and the modified PPtree (implemented in PPtreeExt) are displayed side by side for visual comparison. The user can choose which modifications of PPtree to show. For the other algorithms, only the default parameter choices are used.
Figure 5: Comparing the fits for three-class data with a strong correlation between variables. Modification 1 (different subsetting rules) of PPtree is shown. The modification produces lower error and decision boundaries that are more symmetrically placed between the class clusters.
In Figure 5, three-class data with a strong correlation between variables is simulated. Modification 1 (different subsetting rules) is applied, and rule 1 (Table LABEL:tbl-rules) is used to compute the split point. The modified PPtree achieves an error rate comparable to the original PPtree, with both substantially outperforming rpart. The change in boundaries is subtle, but more satisfying. Between classes S (violet squares) and T (orange triangles), the boundary better matches the separation - it is not pulled away from this split by using LDA on the super-group (T and C).
Figure 6: Comparing the fits for data where one class has two distinct clusters. Modification 2 (multiple splits) of PPtree is used. The modification produces lower error and better decision boundaries.
Figure 6 shows data where one class has two separated clusters, as simulated by Sim-Outlier. The PPtree extension used is modification 2 (multiple splits). It yields a substantial improvement over the other algorithms, both in lower error and more sensible boundaries. The use of multiple splits enables the small group of outliers from class T (orange triangles) to be better classified.
The data in Figure 7 has 5 clusters of different shapes. Modification 2 (multiple splits) is examined. Again, we see that the modified algorithm outperforms PPtree and rpart, in both lower error and better boundaries.
The interactive application provides insight into how modifications to the algorithm affect model fit. Comparing the fitted models follows the principle of visualizing the model in the data space (Wickham et al. 2015). This app has been a valuable tool for debugging and evaluating the algorithmic behavior. By adjusting the simulated data and model parameters, it becomes evident that small changes in the data can result in substantial changes in the fitted model, particularly when projections are involved. By understanding the effect in 2-D, we aim to better understand how the algorithm might also behave in high-dimensional settings.
Figure 7: Comparing the fits for five-cluster Gaussian mixtures. Modification 2 of PPtree is used. This produces lower error and reasonable decision boundaries.
4.2 Numerical performance
We compare the predictive performance for the extended PPtree algorithm with several classification methods, CART, PPtree (LDA, PDA), random forest, PPforest, and SVM, using 94 benchmark data sets. (LDA indicates PP conducted with the linear discriminant index, and PDA indicates PP conducted using the penalized linear discriminant index.) Figure 8 summarizes the performance on data sets where the modified algorithm outperformed other tree methods, and ?@tbl-data-tab2 summarizes the main characteristics of each data set, including the number of observations, predictors, number of classes, class imbalance, and absolute average correlation among predictors. Imbalance quantifies the disparity in class sizes (0 indicates perfectly balanced classes, the biggest proportion size difference between classes), while higher correlation suggests a greater potential for class separation through linear combinations of variables. (The Appendix contains the links to the original data sources, full performance, and data set summaries.)
In the simulation, two-thirds of the observations in each data set are randomly selected for training (using stratified sampling), and the remaining one-third is used for testing and computing the prediction error. This process is repeated 200 times to assess the variance in performance.
Figure 8: Average test error with 95% error bars from 200 training/test splits of benchmark data for four models: rpart, PPtree, PPtreeExt modifications 1 and 2, where the latter performs better than the others. The order of the plots is based on the pattern of errors: at top left PPtreeExt (mod 2) is best but only slightly better than rpart, and at bottom right each of the PPtree versions iteratively performs better than rpart.
Figure 8 summarises the performance for the 18 data sets, where PPtreeExt with modification 2 outperforms the other tree algorithms. In three of these data sets, it also has the lowest error than the ensemble-based classifiers, also. The plots are ordered according to the pattern of test error. In data sets data20-3, data27-a11, data27-a31, data27-b31, data27-c11, data27-c31, data37-5, data38, and data63, only PPtreeExt (mod 2) outperforms rpart. For data34 and data49-2 results are mixed. For the other data sets (data49-3, data22-1, data49-4, data56-1, data67, data69-1, data69-2), each of the methods does better than rpart. The variability across training and test splits is high for some data sets (data38, data22-1). Data sets data20-3, data49-2, and data69-1, representing a range of numerical performance, are chosen to investigate why PPtreeExt performs better, using high-dimensional visualisation in the next section.
Table 2: Overview of benchmark datasets where the proposed tree-based extensions outperform the selected tree methods. The table summarizes the number of observations, predictors, number of classes, class imbalance, and absolute average correlation among predictors. Imbalance quantifies the disparity in class sizes (0 indicates perfectly balanced classes), while higher correlation suggests a greater potential for class separation through linear combinations of variables.
dataset Cases Predictors Groups Imbalance Correlation
data20-3 8143 5 2 0.58 0.45
data27-a11 1747 18 5 0.38 0.24
data27-a31 1834 18 5 0.28 0.27
data27-b31 1424 18 5 0.21 0.24
data27-c11 1111 18 5 0.14 0.27
data27-c31 1448 18 5 0.17 0.27
data37-5 17560 53 5 0.21 0.16
data38 182 12 2 0.43 0.21
data63 6598 166 2 0.69 0.27
data34 1372 4 2 0.11 0.43
data49-2 606 100 2 0.03 1.00
data49-3 606 100 2 0.01 0.99
data22-1 258 5 4 0.25 0.12
data49-4 606 100 2 0.01 0.99
data56-1 7494 16 10 0.01 0.27
data67 20000 16 26 0.00 0.18
data69-1 5000 21 3 0.01 0.30
data69-2 5000 40 3 0.01 0.09
4.3 Visual explanations of performance
Exploring high dimensions is tricky, so it is helpful to set some objectives before leaping into the vast space. We would expect the modified PPtree algorithm to
  • outperform rpart when the best separation between groups is in oblique directions
  • and outperform PPtree when class variances are different, or when there are multiple clusters for any class.
This can be investigated visually using a tour. There are several strategies to use for approaching this:
  1. 1.
    On the data
    1. a.
      Work on two groups at a time, if there are more than 2
    2. b.
      Use a guided tour on the data to find the combinations of the variables where the groups are separated
  2. 2.
    Predict a grid of values that covers the data domain, for the methods interested in comparing, PPtreeExt, rpart, and PPtree, as can be done using the classifly package (Wickham 2022), and then
    1. a.
      subset to the points near the border generated by the classifier.
    2. b.
      record a sequence of projection bases that shows the boundary created by the method in focus.
    3. c.
      use the same projection sequence to view the boundaries for the three different methods, and possibly use a slice tour to focus on the boundary near the center of the data domain.
Figure 9 shows two projections of the data20-3 illustrating why PPtreeExt performs better. The projections are 2-D from the full 5-D space. Color indicates the two classes (0 is blue and 1 is red). The circles represent the projection coefficients. Using a clock dial analogy, in the first projection (a), V3 is pointing towards 1 o’clock, and V4 is pointing towards 10. V2 is also pointing to 1, but is very small. The difference in the two groups is in the direction of 1, so V3 is likely mostly contributing to this, with a small contribution from V2. In the second projection (b) V3 points to 5, V4 points to 2, V5 points to 3, and V2 points to 9. The separation is in the direction of 5 o’clock, suggesting again that V3 is primarily important. Group 0 is split on either side of group 1, which means that the PPtree algorithm would not be able to capture both clusters, but modification 2 would. The separation between the two classes is mostly on V3 but small contributions from other variables help to improve the separation. This means that rpart will not work quite as well because it will focus on splitting V3, ignoring the other variables.
(a) projection 1
(b) projection 2
Figure 9: Two projections from a grand tour of data20-3, from which we can see why the PPtreeExt method is the most effective. Class 0 is split into two clusters on either side of group 1, which means that the original algorithm is not optimal. Secondly, the separation between the two groups is primarily V3 with small contributions from other variables, which gives PPtreeExt a slight edge over rpart.
From Figure 8, for data49-2, we can see that the original PPtree (with LDA) algorithm performs much better than a tree, and the PPtreeExt performs much better again. This data has 100 predictors and 2 classes (Table LABEL:tbl-datatbl), which makes visualising more challenging than data20-3. A useful approach is to pre-process using principal component analysis (PCA) on the predictors, to reduce the number of variables to visualise, while maintaining class structure. In Figure 10 plot (a) shows PC 1 vs PC2, with points coloured by the two classes. PC 1 explains 99.8% of the variability in the predictors, but it explains nothing of the class structure. This is not uncommon for PCA, where an interesting structure can be found in the smaller components. Plot (b) shows PC 2 vs PC 3, and here we see interesting class differences. Each class appears to have three linear pieces pointing in different directions. This supports why PPtreeExt might work well: in the combination of variables, the classes are nearly separable with multiple linear cuts. It’s not feasible to continue working through principal components to tease apart the cluster structure relating to classes, so a good next step is to do PP with the guided tour. The optimal projection is shown in plot (c). It is computed on 26 PCs, and many of these contribute to this projection as seen by the mess of lines inside the circle representing the axes. The two classes are mostly linearly separable in this projection, but there is no clear gap between the clusters. The linear combination with so many variables is what causes some difficulties for a simple tree algorithm. The interesting shapes of the clusters for each class, which appear to be multiple linear pieces, mean that the PPtreeExt can find better separations than PPtree.
(a) PC1 vs PC2
(b) PC2 vs PC3
(c) optimal projection
Figure 10: Examining why data49-2 is best classified using PPtreeExt. Because there are 100 predictors, for visualisation it is useful to reduce the dimension using principal component analysis, and then examine the separation between groups. PC1 explains 99.8% of the overall variability among predictors, but it is not useful for separating the classes (a). PC2 and PC3 (b) have an interesting difference between the two classes, which shows why multiple linear splits might be useful. A guided projection pursuit tour on the first 26 principal components (c) shows that these classes could be separated with a linear combination. This is why the original PPtree classifier does better than the tree algorithm and also why the extension improves the performance.
From Figure 8 we see that all of the PPtree versions perform better than an rpart fit on data69-1. This data has 21 predictors and 3 classes (Table LABEL:tbl-datatbl). This is a sufficiently small-scale problem to examine using a tour, and we initially use a guided tour with the LDA index. This produced the projection shown in plot (a), where we see the primary difference between the three classes. These are each similar size, but oriented in such a way that in this 2-D projection, they form a triangle. Plots (b) and (c) show the boundaries using a slice tour with the same projection as used in (a) for PPtreeExt and rpart, respectively. This uses approach 2 above: created by predicting the class of points randomly generated from a uniform distribution in a 21-D cube of the same range as the data. The slice cuts through to the center of the box. We can see that PPtreeExt essentially fits the data shape neatly, but the rpart boundary is very messy. This data needs oblique cuts, which PPtreeExt provides but rpart does now. The modified algorithm handles the variance differences better than the original PPtree.
(a) data
(b) PPtreeExt
(c) rpart
Figure 11: Examining why data69-1 is best classified using PPtreeExt: (a) best projection from a guided tour shows that the data has three similar linear pieces oriented to form a triangle, (b) boundary produced by the PPtreeExt (mod 2) fitted model, and (c) boundary produced by rpart fitted model. The same projection is used in all plots to allow comparison. The boundary is examined by predicting the class of a uniform random sample of points in the data space, and viewed using a slice tour. The boundary for PPtreeExt (mod 2) mirrors what we learn from the data, that it is carving the data space into three equiangular chunks, but rpart is not capturing these linear combinations.
5 Discussion
This article has presented two modifications to the PPtree algorithm for classification problems implemented in the R package PPtreeExt. These changes result in a more flexible classification method that leverages linear combinations of variables. The performance study showed that the predictive performance of PPtreeExt with modification 2 is better than CART and PPtree in 18 out of 94 data sets. A web app was developed and shared to help understand and debug modifications to the algorithm.
Interpretability is increasingly important because predictions from algorithms may need to be defended in practice. We have shown how to examine the data and the fitted model using high-dimensional visual methods to understand why the PPtree extension performs better than other tree methods on some data sets. This approach is generally useful for assessing and diagnosing model fits.
Ultimately, this may provide better building blocks for a projection pursuit-based classification random forest algorithm (da Silva et al. 2021) as available in the PPforest R package (da Silva et al. 2025). The modified tree algorithm can be incorporated into a projection pursuit forest classifier (da Silva et al. 2025). Given its implicit variable selection mechanism, the ensemble is expected to handle nuisance variables more effectively than a single tree. Using the PDA index (penalised discriminant analysis) also helps when there are many variables. Additionally, the use of bagging should yield improved and more robust performance in problems with nonlinear decision boundaries. This can be accomplished using the current PPforest package.
6 Supplementary Materials
This article was written in Quarto (Allaire et al. 2024), using the R package ggplot2 (Wickham 2016) for visualizations. The files to reproduce the article and results are available at https://github.com/natydasilva/PPtreeExt_paper_JCGS. Animated gifs corresponding to Figure 9, Figure 10 and Figure 11 are in the gifs folder. The code directory contains the scripts for the simulations and to explore the model fits in high dimensions.
The PPtreeExt package provides the modified PPtree algorithm. The Shiny app is available at https://natyds.shinyapps.io/explorapp/ and the app code is provided in the PPtreeExt package, so that it can be used locally for new algorithm developments.